# Chest X-ray Foundation Model with Global and Local Representations Integration

### v1

@@ Abstract
Chest X-ray (CXR) is the most frequently ordered imaging test, supporting diverse clinical tasks from thoracic disease detection to postoperative monitoring. However, task-specific classification models are limited in scope, require costly labeled data, and lack generalizability to out-of-distribution datasets. To address these challenges, we introduce CheXFound, a self-supervised vision foundation model that learns robust CXR representations and generalizes effectively across a wide range of downstream tasks. We pretrain CheXFound on a curated CXR-1M dataset, comprising over one million unique CXRs from publicly available sources. We propose a Global and Local Representations Integration (GLoRI) module for downstream adaptations, by incorporating disease-specific local features with global image features for enhanced performance in multilabel classification. Our experimental results show that CheXFound outperforms state-of-the-art models in classifying 40 disease findings across different prevalence levels on the CXR-LT 24 dataset and exhibits superior label efficiency on downstream tasks with limited training data. Additionally, CheXFound achieved significant improvements on new tasks with out-of-distribution datasets, including opportunistic cardiovascular disease risk estimation and mortality prediction. These results highlight CheXFound’s strong generalization capabilities, enabling diverse adaptations with improved label efficiency. The project source code is publicly available at https://github.com/RPIDIAL/CheXFound.
@@ 1 Introduction
Chest X-ray (CXR) is one of the most commonly-ordered imaging tests worldwide [1]. Clinical CXR interpretation encompasses a broad spectrum of tasks, including detecting diseases associated with the lungs, heart, blood vessel, and bones, as well as monitoring postoperative recovery and the positioning of support devices. With advancements in computer-aided diagnosis, these tasks now extend even further to include opportunistic disease risk assessment, such as cardiovascular disease[2, 3], mortality risk[4], and diabetes[5], among other factors not directly quantifiable by human eyes. Training specialized classification models for each task from scratch poses significant limitations. Such models are typically effective only within a narrow scope of pathologies and struggle to generalize to out-of-distribution datasets. Furthermore, developing these models requires requires extensive labeled datasets, which are both cost-prohibitive and inefficient. These challenges underscore the need for self-supervised models that can learn robust representations and demonstrate superior generalization capabilities across diverse tasks.
@@
Recent advancements in the field of computer vision[6, 7, 8, 9] demonstrate that self-supervised vision models can produce task-agnostic and semantic-rich image representations that achieve improved performance on a broad spectrum of downstream tasks. Such models are called foundation models because of their superior capabilities to adapt to diverse downstream tasks when pretrained on large-scale data. Recent works in self-supervised learning for CXR interpretation adopt a series of advanced training strategies to learn high-quality image representations, including contrastive learning [10], masked image modeling (MIM) [11], and self-distillation [12]. Research further use CXRs and their clinical reports to perform contrastive language-image pretraining [13]. However, these studies have two major limitations. First, they only evaluate model performance for classifying a narrow range of disease findings, without considering the long-tail nature of pathologies in CXR and the opportunistic CXR interpretation tasks, such as cardiovascular disease (CVD) risk estimation and mortality prediction. Second, these studies simply rely on the global image features for disease classification, overlooking the use of CXR representations learned by the foundation models to provide disease-specific local features to enhance performance. Addressing these limitations is pivotal to the development of the CXR foundation models towards clinical applications which often involve interpreting a wide range of disease findings. It also has broader implications by enabling CVD risk estimation and mortality prediction with a routine CXR.
@@
In this work, we introduce CheXFound, a vision foundation model specialized for CXR image analysis that learns high-quality CXR representations and generalizes effectively across a wide range of thoracic disease classification and opportunistic risk estimation tasks. We pretrain CheXFound on a curated CXR-1M dataset, comprising more than one million unique CXRs from 13 publicly available datasets, including MIMIC-CXR [14], CheXpert [15], PadChest [16], CXR14 [17], BRAX [18], VinDr-CXR [19], and CANDID-PTX [20], among others. Our CheXFound model is pretrained via DINOv2 [21], a state-of-the-art self-supervised learning method with strong off-the-shelf linear probe performance. For downstream adaptation, we propose a Global and Local Representations Integration (GLoRI) module. GLoRI is trained on top of the frozen CheXFound model. It uses the attentional principle to compute disease-specific local features and integrates them with the global image features to improve the multilabel classification performance.
@@
We assess CheXFound’s performance on two tiers of CXR interpretation tasks, including thoracic disease classification and opportunistic risk estimation. CheXFound outperforms previous state-of-the-art models such as RAD-DINO [12], EVA-X [11], and CheXzero [13] across 40 disease findings at different prevalence levels on the CXR-LT 24 dataset [22]. Besides, CheXFound demonstrates superior label efficiency, which achieves best-performing results on the Shenzhen, Montgomery, and JSRT datasets with limited training data. We also find that CheXFound achieves significant performance increases compared with its comparisons for the out-of-distribution tasks, including opportunistic CVD risk estimation and mortality prediction on the PLCO dataset [23]. Overall, we demonstrate CheXFound’s strong generalization capabilities across a wide range of downstream tasks on in-distribution and out-of-distribution datasets. CheXFound’s strong representation quality can enable diverse downstream adaptations with improved label efficiency.
@@ 2.2 Foundation Models for Medical Applications
The surge in available data and computational resources have enabled the large-scale pretraining of foundation models. Studies have demonstrated that scaling foundation models in data and model sizes can achieve performance increases across a wide array of downstream tasks [7, 21, 26]. In medical domain, research works have developed multiple categories of foundation models differing in technical approaches and data modalities. Our study is related to vision-centric foundation models. RAD-DINO [12] and EVA-X [11] are two foundation models in CXR domain. Compared to CheXFound with ViT-L pretrained on CXR-1M, these models are limited in model and data scales. Another category of foundation models incorporate vision and text data for multimodal pretraining. CheXzero [13], BiomedCLIP [28], PubMedCLIP [29] use contrastive vision-language pretraining, which is effective in zero-shot classification. Further development of vision-language models takes advantage of instruction-tuning to improve reasoning and detailed description capabilities [30, 31, 32]. Overall, research empirically finds that vision-language models achieve inferior performance than vision-centric foundation models in CXR classification [12]. In this study, we focus on the vision-centric foundation model and investigate its capability for extensitve CXR classification tasks. To the best of our knowledge, our work employs the largest-scale self-supervised pretraining with over 1 million unique CXRs.
@@ 3 Materials and Methods
Table 1: Curation of CXR-1M with publicly available datasets from diverse institutions for self-supervised pretraining. CXR-1M is subsetted into CXR-207K and CXR-744K to evaluate the data scalability of self-supervised models.
@@
CXR14[17] Frontal 14 diseases 112,120
BRAX[18] Frontal, Lateal 14 diseases 40,967
@@
Total number of images in CXR-1M: 1,005,733
3.1 CXR-1M for Pretraining CheXFound
@@
As detailed in Table 1, we curated the CXR-1M dataset for self-supervised pretraining by retrieving in total 1,005,733 unique CXRs from 13 publicly available datasets [14, 15, 16, 17, 36, 18, 37, 38, 19, 35, 34, 20, 33] that were released for various downstream tasks, including disease diagnosis, abnormality detection, foreign objection detection, and segmentation. To learn comprehensive representations for multiview CXR analysis, both frontal-view in PA (posterior-anterior) or AP (anterior-posterior) and lateral-view CXRs were included into CXR-1M.
@@
To evaluate the data scalability of self-supervised pretraining, we further created CXR-207K and CXR-744K, denoting two subsets of CXR-1M, as shown in Table 1. CXR-207K contains approximately 207K CXRs from MIMIC-CXR. CXR-744K contains around 744K CXRs from five datasets: MIMIC-CXR, CheXpert, PadChest, CXR14, and BRAX.
@@
We used DINOv2 [21], a state-of-the-art self-supervised learning method, to pretrain CheXFound on CXR-1M. DINOv2 inherits designs from DINO [8] and iBOT [9] and incorporates two self-distillation objectives: the masked image modeling loss $\mathcal{L}_{\texttt{MIM}}$ and the [CLS] token alignment loss $\mathcal{L}_{\texttt{[CLS]}}$. It uses a teacher-student knowledge distillation architecture as shown in Fig. 1 to learn CXR representations. Masked image modeling uses the teacher network as an online tokenizer, which generates patch tokens from intact images to guide the student network in reconstructing masked patch tokens. This approach enables the student network to learn both visual features and contextual information effectively. On the other hand, the [CLS] token alignment loss $\mathcal{L}_{\texttt{[CLS]}}$ enforces similarity between [CLS] tokens output by the teacher and student networks. This approach aims to train the network to learn high-level class representations with off-the-shelf linear probe capabilities.
3.2 Global and Local Representation Integration for Multilabel Classification
@@
For the downstream evaluation of CheXFound, the linear probe classifier is a pivotal tool to evaluate the quality of pretrained representations. However, the linear probe classifier has limited capability to address the multilabel classification problem commonly seen in CXR interpretation, since it generally relies on the global image features from a single [CLS] token for classifying a wide range of pathologies and lacks essential local details to support the predictions. In contrast, patch tokens from our pretrained CheXFound contain rich CXR representations and high-level contextual information learned via masked image modeling, which can provide disease-specific local features to substantially reduce ambiguities arisen from using the [CLS] token for classifying multiple pathologies. To take advantage of both local and global features for disease classification, we introduce GLoRI (Fig. 2), which utilizes a cross-attention layer with disease queries to summarize patch token features and a skip-connection to integrate the [CLS] token towards final prediction.
@@
To be specific, GLoRI receives output patch tokens $\mathbf{u}^{\texttt{Patch}}\in\mathbb{R}^{N\times D_{\text{model}}}$ from the frozen CheXFound backbone as input, where $D_{\text{model}}$ is the backbone embedding dimension. Since there can be a dimension mismatch between the backbone and GLoRI, we use a linear embedding layer to project $\mathbf{u}^{\texttt{Patch}}$ to the GLoRI dimensional space:
@@
where $\mathbf{u}^{\prime\texttt{Patch}}\in\mathbb{R}^{N\times D_{\text{GLoRI}}}$ is the projected by the linear embedding layer $\text{Linear}^{\text{embed}}(\cdot)$ to the patch token sequence with dimension $D_{\text{GLoRI}}$. In GLoRI, to extract disease-specific local features using the cross-attention layer, we initialize $M$ disease queries corresponding to $M$ disease findings, denoted as $\mathbf{q}\in\mathbb{R}^{M\times D_{\text{key}}}$. The keys $\mathbf{k}\in\mathbb{R}^{M\times D_{\text{key}}}$ and values $\mathbf{q}\in\mathbb{R}^{M\times D_{\text{value}}}$ for the cross-attention layer are from $\mathbf{u}^{\prime\texttt{Patch}}$, which provides rich CXR representations. A scaled dot-product attention module is used to compute attention-pooled features relevant to the query diseases:
@@
$\mathbf{q}^{\prime}=\text{Softmax}\left(\frac{\text{Linear}^{\text{query}}(\mathbf{q})\text{Linear}^{\text{key}}(\mathbf{k})^{T}}{\sqrt{D_{\text{key}}}}\right)\text{Linear}^{\text{value}}(\mathbf{v}),$ (2)

### v2

@@ Abstract
Chest X-ray (CXR) is the most frequently ordered imaging test, supporting diverse clinical tasks from thoracic disease detection to postoperative monitoring. However, task-specific classification models are limited in scope, require costly labeled data, and lack generalizability to out-of-distribution datasets. To address these challenges, we introduce CheXFound, a self-supervised vision foundation model that learns robust CXR representations and generalizes effectively across a wide range of downstream tasks. We pretrained CheXFound on a curated CXR-987K dataset, comprising over approximately 987K unique CXRs from 12 publicly available sources. We propose a Global and Local Representations Integration (GLoRI) head for downstream adaptations, by incorporating fine- and coarse-grained disease-specific local features with global image features for enhanced performance in multilabel classification. Our experimental results showed that CheXFound outperformed state-of-the-art models in classifying 40 disease findings across different prevalence levels on the CXR-LT 24 dataset and exhibited superior label efficiency on downstream tasks with limited training data. Additionally, CheXFound achieved significant improvements on downstream tasks with out-of-distribution datasets, including opportunistic cardiovascular disease risk estimation, mortality prediction, malpositioned tube detection, and anatomical structure segmentation. The above results demonstrate CheXFound’s strong generalization capabilities, which will enable diverse downstream adaptations with improved label efficiency in future applications. The project source code is publicly available at https://github.com/RPIDIAL/CheXFound.
@@ 1 Introduction
Chest X-ray (CXR) is one of the most commonly-ordered imaging tests worldwide [1]. Clinical CXR interpretation encompasses a broad spectrum of tasks, including detecting diseases associated with the lungs, heart, blood vessel, and bones, as well as monitoring postoperative recovery and the positioning of support devices. With advancements in computer-aided diagnosis, these tasks now extend even further to include opportunistic disease risk assessment, such as cardiovascular disease[2, 3], mortality risk[4], and diabetes[5], among other factors not directly quantifiable by human eyes. Training specialized classification models for each task from scratch poses significant limitations. Such models are typically effective only within a narrow scope of pathologies and struggle to generalize to out-of-distribution datasets. Furthermore, developing these models requires extensive labeled datasets, which are both cost-prohibitive and inefficient. These challenges underscore the need for self-supervised models that can learn robust representations and demonstrate superior generalization capabilities across diverse tasks.
@@
Recent advancements in the field of computer vision[6, 7, 8, 9] demonstrate that self-supervised vision models can produce task-agnostic and semantic-rich image representations that achieve improved performance on a broad spectrum of downstream tasks. Such models are called foundation models because of their superior capabilities to adapt to diverse downstream tasks when pretrained on large-scale data. Recent works in self-supervised learning for CXR interpretation adopt a series of advanced training strategies to learn high-quality image representations, including contrastive learning [10], masked image modeling (MIM) [11], and self-distillation [12]. Research further use CXRs and their clinical reports to perform contrastive language-image pretraining [13]. However, these studies have two major limitations. First, most of these proposed models are constrained by either model capacity or pretraining data size, limiting their capability to learn rich and effective representations for long-tail classification, as demonstrated by our detailed experimental results in Fig. 3, showing that existing foundation models achieved degraded performance across all 40 diseases compared to CheXFound with large model capacity and data size. Second, most of these studies simply rely on the global image features for disease detection, overlooking the use of patch embeddings learned by the foundation model to provide disease-specific local features to enhance performance. Addressing these limitations is pivotal to the development of the CXR foundation models towards clinical applications which often involve interpreting a wide range of disease findings. It also has broader implications by enabling CVD risk estimation and mortality prediction with a routine CXR.
@@
In this work, we introduce CheXFound, a vision foundation model specialized for CXR image analysis that learns high-quality CXR representations and generalizes effectively across a wide range of thoracic disease detection and opportunistic risk estimation tasks. We pretrained CheXFound on a curated CXR-987K dataset, comprising approximately 987K unique CXRs from 12 publicly available datasets, including MIMIC-CXR [14], CheXpert [15], PadChest [16], NIH-CXR [17], BRAX [18], and CANDID-PTX [19], among others. Our CheXFound model was pretrained via DINOv2 [20], with specific adaptations in the model initialization strategy and pretraining objective weight to enhance its effectiveness in the CXR domain. For downstream adaptation, we introduce a Global and Local Representations Integration (GLoRI) head that was trained on top of the frozen CheXFound model to perform attention probing. GLoRI employs the attention mechanism to compute fine- and coarse-grained disease-specific local features, which are then integrated with the global image features to improve multilabel classification performance.
@@
We assessed CheXFound’s performance on a wide range of CXR interpretation tasks, from thoracic disease detection, opportunistic risk estimation, and malpositioned tube detection to anatomical structure segmentation. CheXFound outperformed previous state-of-the-art models such as RAD-DINO [12], EVA-X [11], and CheXzero [13] for classifying 40 disease findings at different prevalence levels on the CXR-LT 24 dataset [21]. Besides, CheXFound demonstrated superior label efficiency, achieving best-performing results on the Shenzhen, Montgomery, and JSRT datasets with limited training data. We also found that CheXFound achieved significant performance increases compared with its comparisons for the out-of-distribution tasks, including opportunistic CVD risk estimation and mortality prediction on the PLCO dataset [22], disease detection on the VinDr-CXR dataset [23], malpositioned tube detection on the RANZCR-CLiP dataset [24], and rib segmentation on the VinDr-RibCXR dataset [25]. Overall, we demonstrate CheXFound’s strong generalization capabilities across a wide range of downstream tasks on in-distribution and out-of-distribution datasets. CheXFound’s strong representation quality will enable diverse downstream adaptations with improved label efficiency in future applications.
@@ 2.2 Foundation Models for Medical Applications
Prior to the advent of foundation models, studies used both end-to-end supervised training and supervised fine-tuning methods to develop models for chest X-ray analysis. For example, CheXNet [30] and CheXpert [15] trained convolution neural networks from scratch, achieving better performance than radiologists for the detection of pneumonia and five selected pathologies, respectively. In contrast, supervised fine-tuning methods divided network training into two stages: pretraining to initialize network parameters and fine-tuning to adapt the network to specific tasks. Such methods include CheXtransfer [31], MedAug [32], and MoCo-CXR [33]. CheXtransfer directly fine-tuned a convolutional neural network pretrained on ImageNet for CXR analysis, allowing the model to effectively take advantage of natural domain knowledge to achieve enhanced performance. Both MedAug and MoCo-CXR improved the paradigm of CheXtransfer by pretraining their own models on CXR images via contrastive learning before fine-tuning them on disease detection datasets.
The surge in available data and computational resources have enabled the large-scale pretraining of foundation models. Studies have demonstrated that scaling foundation models in data and model sizes can achieve performance increases across a wide array of downstream tasks [7, 20, 28]. In medical domain, studies have developed multiple categories of foundation models that use different pretraining approaches and data modalities. Our study belongs to the family of vision-centric foundation models. In the domain of CXR image analysis, several studies have explored various strategies for developing vision-centric foundation models, including RAD-DINO [12], RayDINO [34], EVA-X [11], Medical MAE [35], Foundation Ark [36], and Adam [37]. However, as summarized in Table 1, most of these related studies remain limited in either model capacity or data scale. Models such as [12, 11, 35, 36, 37] typically employed the ViT-Base or Swin Transformer-Base architecture, with up to 88M parameters, and were pretrained on up to 838K CXRs from five public datasets. To further investigate the scalability, this work scales CheXFound to the ViT-Large architecture and pretrains it on 987K CXRs from 12 different sources. Among the above CXR foundation models, only RayDINO was based on the ViT-Large architecture, as adopted in this work. However, RayDINO did not sufficiently exploit the benefit of data diversity, relying on only four commonly used public datasets for pretraining. In contrast, our work expanded the pretraining dataset with eight additional sources, resulting in 124K more pretraining images.
Table 1: Summary of the network architectures and pretraining datasets utilized by recent vision-centric foundation models.
Models Architectures # Parameters # CXRs # Data sources Medical MAE [35] ViT-Base 86M 509K 3 public datasets Adam [37] ResNet-50 25M 224K 1 public datasets Foundation Ark [36] Swin Transformer-Base 88M 704K 4 public datasets EVA-X [11] ViT-Base 86M 520K 3 public datasets RAD-DINO [12] ViT-Base 86M 838K 5 public datasets + private dataset RayDINO [34] ViT-Large 307M 863K 4 public datasets CheXFound ViT-Large 307M 987K 12 public datasets
Another category of foundation models incorporate vision and text data for multimodal pretraining. CheXzero [13], BiomedCLIP [38], PubMedCLIP [39] use contrastive vision-language pretraining, which is effective in zero-shot classification. Further development of vision-language models takes advantage of instruction-tuning to improve reasoning and detailed description capabilities [40, 41, 42]. Overall, research empirically finds that vision-language models achieve inferior performance than vision-centric foundation models in CXR classification [12]. In this study, we focus on the vision-centric foundation model and investigate its capability for extensive CXR classification tasks. To the best of our knowledge, our work employs the largest-scale self-supervised pretraining with over 1 million unique CXRs.
@@ 3 Materials and Methods
Table 2: Curation of CXR-987K with publicly available datasets from diverse institutions for self-supervised pretraining. CXR-987K is subsetted into CXR-207K and CXR-744K to evaluate the data scalability of self-supervised models.
@@
NIH-CXR[17] Frontal 14 diseases 112,120
BRAX[18] Frontal, Laterall 14 diseases 40,967
@@
Total number of images in CXR-987K: 987,733
3.1 Pretraining CheXFound with Large-scale Dataset
@@
As detailed in Table 2, we curated the CXR-987K dataset for self-supervised pretraining by retrieving in total 987,733 unique CXRs from 12 publicly available datasets [14, 15, 16, 17, 46, 18, 47, 48, 23, 45, 44, 19, 43] that were released for various downstream tasks, including disease diagnosis, abnormality detection, foreign objection detection, and segmentation. To learn comprehensive representations for multiview CXR analysis, both frontal-view in PA (posterior-anterior) or AP (anterior-posterior) and lateral-view CXRs were included into CXR-987K.
@@
To evaluate the data scalability of self-supervised pretraining, we further created CXR-207K and CXR-744K, denoting two subsets of CXR-987K, as shown in Table 2. CXR-207K contains approximately 207K CXRs from MIMIC-CXR. CXR-744K contains around 744K CXRs from five datasets: MIMIC-CXR, CheXpert, PadChest, NIH-CXR, and BRAX.
@@
We used DINOv2 [20], a state-of-the-art self-supervised learning method, to pretrain CheXFound on CXR-987K. DINOv2 inherits designs from DINO [8] and iBOT [9] and incorporates two self-distillation objectives: the masked image modeling loss $\mathcal{L}_{\texttt{MIM}}$ and the [CLS] token alignment loss $\mathcal{L}_{\texttt{[CLS]}}$. It uses a teacher-student knowledge distillation architecture as shown in Fig. 1 to learn CXR representations. Masked image modeling uses the teacher network as an online tokenizer, which generates patch tokens from intact images to guide the student network in reconstructing masked patch tokens. This approach enables the student network to learn both visual features and contextual information effectively. On the other hand, the [CLS] token alignment loss $\mathcal{L}_{\texttt{[CLS]}}$ enforces similarity between [CLS] tokens output by the teacher and student networks. This approach aims to train the network to learn high-level class representations with off-the-shelf linear probe capabilities.
3.2 Global and Local Representations Integration for Multilabel Classification
@@
For the downstream evaluation of CheXFound, the linear probe classifier is a pivotal tool to evaluate the quality of pretrained representations. However, the linear probe classifier has limited capability to address the multilabel classification problem commonly seen in CXR interpretation, since it generally relies on the global image features from a single [CLS] token for classifying a wide range of pathologies and lacks essential local details to support the predictions. In contrast, patch tokens from our pretrained CheXFound contain rich CXR representations and high-level contextual information learned via masked image modeling, which can provide disease-specific local features to substantially reduce ambiguities arisen from using the [CLS] token for classifying multiple pathologies. To take advantage of both local and global features for multilabel disease classification, we introduce GLoRI (Fig. 2), which consists of two local-feature branches that use an adaptive temperature module to provide fine-grained local features and a pyramid patch merging module to provide coarse-grained local features respectively and a skip connection to integrate the global image features from the [CLS] token for multilabel disease classification.
3.2.1 Encoding Fine-grained Local Features with Adaptive Temperatures
@@
GLoRI receives output patch tokens $\mathbf{u}^{\texttt{Patch}}\in\mathbb{R}^{N\times D_{\text{model}}}$ from the frozen CheXFound backbone as input, where $D_{\text{model}}$ is the backbone embedding dimension. Since there can be a dimension mismatch between the backbone and GLoRI, we use a linear embedding layer to project $\mathbf{u}^{\texttt{Patch}}$ to the GLoRI dimensional space:
@@
where $\mathbf{u}^{\prime\texttt{Patch}}\in\mathbb{R}^{N\times D_{\text{GLoRI}}}$ is the projected by the linear embedding layer $\text{Linear}^{\text{embed}}(\cdot)$ to the patch token sequence with dimension $D_{\text{GLoRI}}$. In GLoRI, to extract disease-specific local features using the multi-head attention layer, we initialize $M$ disease queries corresponding to $M$ disease findings, denoted as $\mathbf{q}\in\mathbb{R}^{M\times D_{\text{key}}}$. The keys $\mathbf{k}\in\mathbb{R}^{N\times D_{\text{key}}}$ and values $\mathbf{v}\in\mathbb{R}^{N\times D_{\text{value}}}$ for the multi-head attention layer are derived from $\mathbf{u}^{\prime\texttt{Patch}}$ to provide rich CXR representations.
Additionally, to improve GLoRI’s capability to focus on small or diffuse abnormal regions, we introduce an adaptive temperature module to adjust the attention in the multi-head attention layer. The adaptive temperature module projects the average-pooled features of $\mathbf{u}^{\texttt{Patch}}$ through a multilayer perceptron with a tanh activation function and then applies a natural exponential function to generate the temperature vector $\boldsymbol{\uptau}\in\mathbb{R}^{M}$, where $M$ corresponds to the number of disease queries. We incorporate $\boldsymbol{\uptau}$ into the scaled dot-product attention mechanism to capture fine-grained local features:
@@
${\color[rgb]{0,0,0}\mathbf{q}^{\prime}=\text{Softmax}\left(\frac{\mathbf{q}\mathbf{k}^{T}}{\sqrt{D_{\text{key}}}\boldsymbol{\uptau}}\right)\mathbf{v},}$ (2)
where $\mathbf{q}^{\prime}\in\mathbb{R}^{M\times D_{\text{key}}}$ denotes the output disease queries that capture fine-grained local features. $\boldsymbol{\uptau}$ adjusts the distribution of attention weights for each disease query to be either sharper or smoother with a value lower or larger than 1 respectively.
3.2.2 Encoding Coarse-grained Local Features with Pyramid Patch Merging
To effectively summarize contextual information, we incorporate a pyramid patch merging module to integrate multi-scale patch embeddings at low resolutions to provide coarse-grained local features. The pyramid patch merging module merges $8\times 8$, $4\times 4$, and $2\times 2$ adjacent patch tokens in $\mathbf{u}^{\texttt{Patch}}$ via average pooling to generate multi-scale patch embeddings. Then, it projects these patch embeddings to a dimension of $\frac{D_{\text{model}}}{3}$ via a linear layer and a ReLU activation, upsamples them to the original resolution, and concatenate them to construct patch embeddings with a dimension of $D_{\text{model}}$. We then apply a layer normalization to the patch embeddings before inputting them to the multi-head attention to encode coarse-grained local features.
3.2.3 Integrating Global Image Features